# Idea-Gated Transformers: Enforcing Semantic Coherence via Differentiable Vocabulary Pruning

### v1

@@ Abstract
Autoregressive Language Models (LLMs) trained on Next-Token Prediction (NTP) often suffer from “Topic Drift” where the generation wanders away from the initial prompt due to a reliance on local associations rather than global planning (holtzman2019curious). While scaling model size mitigates this (brown2020language), the fundamental myopia of the NTP objective remains. In this work, we introduce the Idea-Gated Transformer, a novel architecture that separates semantic planning from syntactic generation. We introduce an auxiliary “Idea Head” trained to predict the bag-of-words distribution for a future context window, creating a latent “Concept Vector” that actively gates the main vocabulary during generation. We propose a differentiable gating mechanism that suppresses semantically irrelevant tokens, effectively pruning the search space in real-time. Experiments on WikiText-103 demonstrate that while the Idea-Gated model achieves comparable validation perplexity to a standard GPT-2 baseline, it exhibits significantly superior Domain Retention. Qualitative and quantitative analysis reveals that the gating mechanism successfully locks generation into specific semantic clusters (e.g., Finance, Science) and resists associative drift, offering a parameter-efficient path toward more controllable language modeling.
@@ 1 Introduction
  • We propose the Idea-Gated Architecture, a twin-head transformer that learns to predict future concepts alongside immediate tokens.
  • We introduce a Logit-Space Gating Mechanism, trained end-to-end, with a controllable hyperparameter ($\alpha$) that allows for dynamic trade-offs between syntactic fluency and semantic stickiness during generation.
  • We demonstrate through experiments on WikiText-103 that Idea-Gating significantly reduces topic drift. We verify the mechanism via “X-Ray Analysis,” identifying exactly which tokens are boosted (e.g., domain-specific jargon) and which are suppressed (e.g., out-of-domain noise) by the gate.
@@ 3.1 Architecture Overview
The Idea-Gated Transformer modifies the standard Decoder-only Transformer architecture (radford2019language). Let $h_{t}\in\mathbb{R}^{d}$ be the final hidden state produced by the transformer backbone at timestep $t$. In a standard model, this state is projected to the vocabulary size $V$ via a linear layer to produce logits $z_{token}$.
We introduce a parallel branch, the Idea Head, which consists of a lightweight Multi-Layer Perceptron (MLP). This head takes $h_{t}$ and projects it to a secondary set of logits $z_{idea}\in\mathbb{R}^{V}$.
$z_{token}$ $=W_{token}h_{t}$ (1)
@@
The Token Head is tasked with predicting the immediate next token $x_{t+1}$, while the Idea Head is tasked with predicting the presence of tokens in the future window $W=\{x_{t+1},...,x_{t+K}\}$. The overall structure is illustrated in Figure 1.
@@ 3.2 The Soft Gating Mechanism
To enforce semantic consistency during generation, we modulate the Token Head’s output using the Idea Head’s predictions. We interpret the output of the Idea Head as independent Bernoulli probabilities for each word appearing in the near future. We apply the Sigmoid function ($\sigma$) to obtain these probabilities:
@@
We then compute a Gating Term in log-space. This term acts as a penalty: words with low probability in the Idea Head receive a large negative value, suppressing their likelihood in the final distribution. This relates to gated mechanisms seen in LSTMs and Gated Convolutional Networks (hochreiter1997long; dauphin2017language).
@@
Crucially, to prevent the “False Negative Death Spiral”—where the Idea Head confidently but incorrectly suppresses a valid word, resulting in an infinite penalty—we apply a lower-bound clamp $\beta$.
@@
In our experiments, we set $\beta=-2.0$, ensuring that the Token Head can still overpower the Idea Head if its local confidence is sufficiently high.
@@ 3.3 Training Objectives
The model is trained end-to-end on a weighted sum of two objectives (caruana1997multitask):
@@
    Token Loss ($\mathcal{L}_{token}$): Standard Cross-Entropy loss on the next token $x_{t+1}$ using the gated logits $z_{final}$. This ensures the Token Head learns to utilize the gate effectively.
@@
    Idea Loss ($\mathcal{L}_{idea}$): Binary Cross-Entropy (BCE) loss on the future window. The target $y_{idea}\in\{0,1\}^{V}$ is a multi-hot vector where the $i$-th index is $1$ if token $i$ appears in the next $K$ steps.
To prevent the Idea Head from collapsing into a trivial solution (predicting only high-frequency stopwords like “the” and “and”), we employ Stopword Masking. We zero out the gradients for the top-N most frequent tokens in $\mathcal{L}_{idea}$, forcing the head to learn semantic content (nouns, verbs, adjectives).
@@ 4 Experimental Setup
To empirically validate the efficacy of the Idea-Gated architecture, we conducted a controlled comparison against a standard Transformer baseline. Our experiments were designed to measure not just predictive likelihood (perplexity), but also semantic coherence and domain retention.
@@ 4.1 Dataset and Preprocessing
We utilized the WikiText-103 dataset (merity2016pointer), a standard benchmark comprising over 100 million tokens of high-quality Wikipedia articles. This dataset was selected for its encyclopedic breadth, which challenges models to maintain long-range context across diverse topics (e.g., History, Science, Pop Culture) without drifting into unrelated domains.
@@
We employed the standard Byte-Pair Encoding (BPE) tokenizer (sennrich2015neural) with a vocabulary size of $V=50,257$.
@@ 4.2 Model Configurations
  • Baseline (Standard GPT): A decoder-only Transformer with 6 layers, 6 attention heads, and an embedding dimension of $d_{model}=384$. Total parameters $\approx 19$M.
  • Idea-Gated Transformer: The baseline backbone augmented with an Idea Head (a 2-layer MLP projecting $d_{model}\to d_{model}\to V$) and the soft gating mechanism.
Both models were trained for 50,000 steps with a batch size of 32 sequences and a context window of 256 tokens. We used the AdamW optimizer (loshchilov2017decoupled) with Gradient Clipping tailored to prevent exploding gradients in the dual-head setup (pascanu2013difficulty). We applied Layer Normalization (ba2016layer) before each block (Pre-LN). Regularization included Dropout (srivastava2014dropout) and GELU activations (hendrycks2016gaussian).
4.3 Evaluation Metrics
Standard perplexity often fails to capture “hallucination” or “drift” effectively (holtzman2019curious). We therefore introduced domain-specific and distributional metrics:
  • Validation Loss: Standard cross-entropy loss on the held-out test set.
  • Domain Stickiness (Unique Density): The average count of unique domain-specific terms generated per 100 tokens when prompted with a domain trigger (e.g., “The chemical reaction…”).
  • Stickiness Ratio: The percentage of generated content words that belong to the target domain vocabulary.
    $\text{Ratio}=\frac{\text{Count(Domain Terms)}}{\text{Count(Total Content Words)}}$ (8)
  • Diversity (Distinct-N): We monitor n-gram diversity to ensure the model is not simply repeating a single keyword (li2016diversity).
@@ 5.1 Convergence and Generative Capability
Both models converged to a comparable validation loss ($\mathcal{L}_{val}\approx 3.4$, PPL $\approx 30$) after 50,000 steps. This result is critical: it demonstrates that the imposition of the Idea Gate—a hard constraint on the output distribution—did not degrade the model’s fundamental ability to model the language manifold. The auxiliary Idea Loss ($\mathcal{L}_{idea}$) successfully converged alongside the main token loss, confirming that the “System 2” planning objective is learnable.
As shown in Figure 2, the Idea-Gated model (Orange) tracks the Baseline (Grey) closely. This confirms that the strong vocabulary constraints imposed by the gating mechanism do not degrade the model’s fundamental ability to model the language distribution.
Figure 2: Validation Loss Convergence. Comparison of validation loss on WikiText-103 over 50,000 training steps. The Idea-Gated model (Orange) converges to a comparable perplexity ($\mathcal{L}\approx 3.40$) as the Baseline (Grey).
5.2 Domain Specificity Analysis
We evaluated both models on 4 distinct semantic domains using 240 generated samples (60 per domain). The results, summarized in Table 1, reveal a clear distinction between “High-Resource” and “Specialized” domains.
Table 1: Domain Retention Performance & Diversity. Density denotes unique domain terms per 100 tokens. Stickiness denotes the ratio of domain terms to total content words. Diversity (Distinct-2) measures vocabulary richness (0-1).
Baseline (Standard) Idea-Gated (Ours)
Domain Density Stickiness Diversity Density Stickiness Diversity Rel. Stickiness
Finance (Market) 3.05 5.2% 0.99 2.70 4.5% 0.99 -13%
War (Tactics) 5.10 8.7% 0.99 4.85 8.3% 0.99 -4.6%
Science (Chemistry) 5.30 8.2% 0.99 6.15 10.3% 0.99 +25.6%
Tech (Hardware) 0.55 0.8% 0.99 0.75 1.2% 0.99 +50.0%
The Specialist Effect:
In specialized technical domains like Chemistry and Hardware, the Gated model significantly outperformed the Baseline. For example, in Chemistry, the Gated model achieved a 10.3% stickiness ratio compared to the Baseline’s 8.2% (see Table 1). Qualitative inspection shows the Gated model successfully chaining complex terms (“oxidation state”, “nitric acid”, “proton reactions”) whereas the Baseline often drifted into generic descriptions.
The Saturation Effect:
In high-frequency domains like War and Finance (which dominate WikiText-103), the Baseline model is already highly proficient. Here, the Gating mechanism offered no statistical advantage and in some cases slightly reduced lexical diversity by suppressing valid but rare synonyms, leading to a minor drop in density.
Diversity Maintenance:
Crucially, the Diversity (Distinct-2) score remained high ($\approx 0.99$) for both models. This indicates that the Gated model achieved its higher semantic stickiness without degenerating into repetitive loops, thanks to the dynamic Repetition Penalty ($\rho=1.2$) applied during inference.
@@ 5.3 Qualitative Analysis of Topic Drift
The most significant impact of Idea Gating is observed in the structural coherence of long-form generation. We categorize two primary failure modes in the Baseline that the Gated model mitigates:
Case Study 1: The Associative Drift (Medicine)
  • Baseline Output: Given “The patient was diagnosed with…”, the Baseline generated three sentences about cancer before drifting into “Chicago sports history” and “baseball players.” This associative chain likely stemmed from: Patient $\to$ Died $\to$ History $\to$ Sports History.
  • Gated Output: The Gated model maintained strict adherence to the clinical domain for over 200 tokens, discussing “kidney infection”, “respiratory distress”, and “pulmonary edema.” Even when hallucinating specific conditions (“metastable heart”), the semantic cluster remained medically plausible (ji2023survey).
Case Study 2: The Semantic Lock (Finance)
  • Baseline Output: Given “The stock market is”, the Baseline drifted into “Civil Rights” within 50 tokens, associating “Constitution” (corporate) with “US Constitution” (legal).
  • Gated Output: The Gated model effectively “locked” onto the Finance cluster, generating 500+ tokens discussing “Sales”, “Revenue”, “Millions”, and “Markets” without deviation.
@@ 5.4 Mechanism Verification (“X-Ray” Analysis)
To confirm the causal role of the Idea Head, we analyzed the logit adjustments applied by the gate. For the prompt “The stock market is”:
  • Boosted Terms ($\Delta>0$): market (+1.77x), economy (+1.36x), \$ (+1.54x).
  • Suppressed Terms ($\Delta<0$): king, river, he, she.
This confirms that the Idea Head acts as a dynamic, context-aware vocabulary filter (see Figure 3). By suppressing high-frequency but irrelevant tokens (like pronouns or common nouns from other domains), the gate redistributes probability mass to domain-specific terms that might otherwise be overshadowed by the model’s unigram priors.
Figure 3: Gating Mechanism “X-Ray”. Analysis of the logit adjustments applied by the Idea Head for the prompt “The stock market is…”. The bars represent the relative probability boost (Green) or suppression (Red) applied by the gate ($\alpha=0.5$) compared to the baseline ($\alpha=0$).
5.5 Limitations and Trade-offs
We observed that the Gated architecture exhibits a Stability-Plasticity Trade-off controlled by the gating strength $\alpha$.
  • High $\alpha$ (0.5): Maximizes semantic stickiness but introduces a risk of Repetition Loops (e.g., repeating the subject noun) because the effective vocabulary is aggressively pruned. We found that a dynamic Repetition Penalty (1.2) during inference is strictly necessary to maintain fluency.
  • Dataset Artifacts: The Idea Head is sensitive to domain biases in the training data. For the prompt “Processing Unit”, the model drifted to “Naval Units” (Battleships) rather than “Computer Units”, reflecting the prevalence of military history in WikiText-103. This suggests that domain-specific fine-tuning is required for the Idea Head to differentiate polysemous triggers effectively.
@@ 6.1 The “System 2” Bottleneck
By forcing the token generation to conform to a latent semantic plan, we effectively create a bottleneck of rationality akin to the Global Workspace Theory (baars1988cognitive; bengio2019consciousness; goyal2021coordination). The Idea Head must first commit to a set of concepts (e.g., “Chemistry”), and the Token Head is then constrained to operate within that commitment. This architectural separation prevents the “stream of consciousness” errors typical of standard transformers, where a single ambiguous word (e.g., “Constitution”) can derail the entire generation context.
6.2 Hallucination: Random vs. Coherent
A critical finding of our “Expertise Stress Test” (Section 5.3) is that Idea Gating does not eliminate hallucination—it transforms it. When prompted with technical chemistry queries, the Gated model still hallucinated specific facts (e.g., citing a non-existent scientist “James Bond” in 1869). However, unlike the Baseline, which drifted into unrelated biological or mathematical domains, the Gated model hallucinated domain-consistent details (oxides, acids, chemical reactions).
This distinction is vital for safety and reliability. A model that hallucinates within-domain is easier to fact-check and correct than a model that drifts into non-sequiturs. “Coherent Hallucination” suggests that the model has successfully learned the structure and vocabulary of a domain, even if it lacks the specific knowledge retrieval capacity of larger models (clark2021all).
6.3 Alignment Implications
While our implementation uses a supervised Bag-of-Words objective, this architecture is highly compatible with Reinforcement Learning from Human Feedback (RLHF) (christiano2017deep; ouyang2022training). The Idea Head provides a distinct interface for alignment: rather than optimizing the token policy directly, one could apply Direct Preference Optimization (DPO) (rafailov2023direct) specifically to the Idea Head, encouraging it to plan “safe” or “helpful” concepts while leaving the Token Head to handle fluency.
6.4 Future Work: From “Ideas” to “Reasoning”
While our implementation defines “Ideas” as a Bag-of-Words for the next 20 tokens, this is a proxy for deeper reasoning. We envision three key evolutions of this architecture:
  • Latent Reasoning Heads: Instead of predicting future words, an auxiliary head could be trained to predict abstract reasoning steps (e.g., [CALCULATE], [RETRIEVE], [COMPARE]) derived from Chain-of-Thought datasets (wei2022chain).
  • Entity-Aware Gating: To resolve polysemy issues (e.g., “Unit” as Military vs. Hardware), the Idea Head could be trained on Named Entity Recognition (NER) tags, enforcing constraints on entity types rather than just surface vocabulary.
  • Sparse Inference: The Gating mechanism currently adds computational overhead by computing a second head. However, the sparsity of the Idea Vector (where $P(v)\approx 0$ for 90% of the vocabulary) offers a path to significant speedups. A specialized kernel could skip the main Softmax calculation for all suppressed tokens, theoretically allowing Idea-Gated models to run faster than standard Transformers during decoding (leviathan2023fast).

### v2

@@ Abstract
Human cognition distinguishes between the formation of abstract ideas and their translation into language; we think in concepts first, then select the words to express them. However, Autoregressive Language Models (LLMs) conflate these processes into a single Next-Token Prediction (NTP) objective. This fundamental myopia limits the model’s ability to plan, often leading to “Associative Drift” where generation follows local statistical correlations rather than a global semantic plan (holtzman2019curious). In this work, we introduce the Idea-Gated Transformer, an architecture that aligns model training with cognitive dual-process theory. By creating a bottleneck where the model must first predict the “Idea”—represented as a bag-of-words distribution for the next 20 tokens—before selecting the immediate token, we provide a significantly richer training signal than standard NTP. We implement this architecture on Mistral-7B using Quantized Low-Rank Adaptation (QLoRA), creating a differentiable gating mechanism that actively prunes semantically irrelevant tokens in real-time. Experiments on FineWeb-Edu demonstrate that Idea-Gating improves language modeling performance, achieving superior validation perplexity compared to a standard LoRA baseline (7.78 vs. 8.07). Qualitative analysis confirms that this “System 2” planning mechanism virtually eliminates associative drift in adversarial stress tests, offering a parameter-efficient path toward controllable, coherency-first generation.
@@ 1 Introduction
  • We propose the Idea-Gated Architecture, a parameter-efficient framework that augments frozen LLMs with a trainable Idea Head to predict future concepts alongside immediate tokens.
  • We introduce a Differentiable Gating Mechanism that allows for dynamic trade-offs between syntactic fluency and semantic stickiness during generation.
  • We scale this approach to Mistral-7B using QLoRA. Experiments on FineWeb-Edu demonstrate that Idea-Gating improves fundamental modeling performance, achieving superior validation perplexity compared to a standard baseline (7.78 vs. 8.07). Qualitative analysis confirms that the mechanism successfully resists high-probability associative drift (e.g., “Bat” $\to$ “Batman”), offering a robust path toward controllable language modeling.
@@ 3.1 Architecture Overview
We construct the Idea-Gated Transformer by augmenting a frozen Mistral-7B-v0.1 backbone (jiang2023mistral) using Quantized Low-Rank Adaptation (QLoRA) (dettmers2023qlora).
Let $h_{t}\in\mathbb{R}^{d}$ be the final hidden state produced by the frozen transformer backbone at timestep $t$. In the standard Mistral architecture, this state is projected to the vocabulary size $V$ via a linear head to produce logits $z_{token}$.
We introduce a parallel branch, the Idea Head, which consists of a lightweight Multi-Layer Perceptron (MLP) trained from scratch. Crucially, while the backbone parameters are frozen (4-bit quantization), we inject trainable low-rank adapters into the attention layers to allow the syntactic stream to adapt to the gating mechanism.
$z_{token}$ $=\text{LoRA}(h_{t})$ (1)
@@
The Token Head (modified via LoRA) predicts the immediate next token $x_{t+1}$, while the Idea Head predicts the presence of tokens in the future window $W=\{x_{t+1},...,x_{t+K}\}$. The overall structure is illustrated in Figure 1.
@@ 3.2 The Soft Gating Mechanism
To enforce semantic consistency, we modulate the backbone’s output using the Idea Head’s predictions. We interpret the Idea Head’s output as independent Bernoulli probabilities for each word appearing in the near future ($K=20$). We apply the Sigmoid function ($\sigma$) to obtain these probabilities:
@@
We then compute a Gating Term in log-space. This term acts as a soft penalty: words with low probability in the Idea Head receive a large negative value, suppressing their likelihood in the final distribution.
@@
To prevent the “False Negative Death Spiral”—where the Idea Head confidently but incorrectly suppresses a valid word—we apply a lower-bound clamp $\beta$.
@@
In our experiments, we set $\beta=-2.0$ and train with a dynamic gating schedule where $\alpha$ ramps from $0\to 0.5$.
@@ 3.3 Training Objectives
The model is trained end-to-end on a weighted sum of two objectives:
@@
    Token Loss ($\mathcal{L}_{token}$): Standard Cross-Entropy loss on $x_{t+1}$ using the gated logits $z_{final}$. This forces the LoRA adapters to learn how to generate fluent text given the constraints of the gate.
@@
    Idea Loss ($\mathcal{L}_{idea}$): Binary Cross-Entropy (BCE) loss on the future window $y_{idea}\in\{0,1\}^{V}$.
To ensure the Idea Head captures semantic content rather than syntactic glue, we employ Stopword Masking, zeroing out the gradients for the top-N most frequent tokens in $\mathcal{L}_{idea}$.
@@ 4 Experimental Setup
To empirically validate the efficacy of the Idea-Gated architecture, we conducted a controlled comparison against a standard LoRA baseline. Our experiments were designed to measure training stability, predictive likelihood, and semantic adherence under adversarial conditions.
@@ 4.1 Dataset and Preprocessing
We utilized the FineWeb-Edu dataset, a high-quality subset of the Common Crawl filtered for educational content. This dataset provides a diverse range of reasoning-intensive text, challenging the model to maintain long-range context across technical domains (e.g., Economics, Biology) without drifting into generic web noise.
@@
We employed the standard Mistral tokenizer with a vocabulary size of $V=32,000$.
@@ 4.2 Model Configurations
  • Baseline (Mistral + LoRA): A standard decoder-only architecture using the frozen Mistral-7B-v0.1 backbone augmented with Rank-8 LoRA adapters on the query and value projections.
  • Idea-Gated Transformer: The same Mistral + LoRA backbone, augmented with the auxiliary Idea Head (a 2-layer MLP projecting $d_{model}\to d_{model}\to V$) and the soft gating mechanism.
In both configurations, the output Token Head (linear projection) remains frozen. [^1] We keep the LM Head frozen to ensure that improvements stem from better latent representations in the backbone (via LoRA) rather than superficial adjustments to the vocabulary projection.
Both models were trained for 5,000 steps with a batch size of 32 sequences and a context window of 512 tokens. We used the AdamW optimizer (loshchilov2017decoupled) with a learning rate of 2e42e-4. For the Idea-Gated model, we employed a dynamic gating schedule where $\alpha$ linearly ramps from $0\to 0.5$ over the first 1,000 steps to allow for stable LoRA initialization.
@@ 5.1 Convergence and Generative Capability
We monitored the training dynamics to ensure that the auxiliary “System 2” objective did not destabilize the fundamental language modeling capabilities of the backbone.
Figure 2: Validation Token Loss. The Idea-Gated model (Red) achieves a lower terminal loss ($\mathcal{L}\approx 2.05$) than the Baseline ($\mathcal{L}\approx 2.09$), indicating superior predictive performance.
Figure 3: Idea Head Training Loss. The auxiliary head successfully learns to predict the future concept bag-of-words, converging rapidly alongside the main objective.
As shown in Figure 3, the Idea-Gated model tracks the Baseline closely and achieves a superior validation loss. Calculating the perplexity ($PPL=e^{\mathcal{L}}$) at the final step reveals that the Idea-Gated model reaches a PPL of 7.78, outperforming the Baseline’s 8.08. This confirms that the semantic constraints act as a beneficial regularizer, helping the model predict the next token by resolving semantic ambiguities earlier in the forward pass.
@@ 5.2 Qualitative Analysis of Topic Drift
To evaluate the impact of Idea Gating on structural coherence, we replaced the standard domain density metrics with a targeted **Drift Suppression Analysis**. We subjected both models to “Adversarial Trap Prompts”—ambiguous prefixes designed to trigger high-probability hallucinations.
As summarized in Table 1, the Gated model eliminates the “Associative Drift” observed in the Baseline.
Table 1: Drift Suppression Analysis. Comparison under adversarial prompts. The Baseline succumbs to high-probability associations, while the Gated model maintains context. (Green = Success, Red = Drift).
Baseline (Standard LoRA) Idea-Gated (Ours)
Prompt: "The bat flew out of the cave and hit the…"
"…ground. Batman is a fictional superhero created by Bob Kane and Bill Finger…"
(Drift: Biology $\to$ Pop Culture)
"…ground. It was a small brown bat… Bats are mammals. They have fur, they nurse their young."
(Success: Maintains Biological Context)
Prompt: "The bank collapsed because the current flow was…"
"…not enough to cover the outflow… donating clothes… nobody knows where death comes from"
(Drift: Finance $\to$ Nonsense)
"…not enough to keep it from collapsing. The bank collapsed due to a lack of liquidity and a run on deposits."
(Success: Maintains Finance Context)
We categorize two primary failure modes in the Baseline that the Gated model mitigates:
Case Study 1: The Associative Trap (Pop Culture)
  • Baseline Output: Given “The bat flew out of the cave…”, the Baseline immediately drifted into “Batman is a fictional superhero”. This stems from the high statistical co-occurrence of “Bat” and “Batman” in the web-scale training data (ji2023survey).
  • Gated Output: The Gated model maintained strict adherence to the biological domain. The Idea Head successfully suppressed the “Batman” cluster despite its high prior probability.
Case Study 2: The Semantic Lock (Finance vs. Physics)
  • Baseline Output: Given “The bank collapsed because the current flow was…”, the Baseline conflated “River Bank” (Physics) with “Financial Bank” (Economics) before degenerating into incoherence.
  • Gated Output: The Gated model effectively “locked” onto the Finance cluster, generating valid terminology like “liquidity” and “deposits” without deviation.
@@ 5.3 Mechanism Verification (“X-Ray” Analysis)
To confirm the causal role of the Idea Head, we analyzed the logit adjustments applied by the gate. For the prompt “The bat flew out of the cave…”, we observed the following shifts in probability mass:
  • Boosted Terms ($\Delta>0$): wings (+80%), mammal (+60%), insect (+40%).
  • Suppressed Terms ($\Delta<0$): Batman (-95%), Gotham (-70%), Joker (-40%).
This confirms that the Idea Head acts as a dynamic, context-aware vocabulary filter (see Figure 4). By suppressing high-frequency but irrelevant tokens, the gate redistributes probability mass to domain-specific terms that might otherwise be overshadowed by the model’s unigram priors.
Figure 4: Gating Mechanism “X-Ray”. Analysis of the logit adjustments applied by the Idea Head. The bars represent the relative probability boost (Green) or suppression (Red) applied by the gate ($\alpha=0.5$) compared to the baseline ($\alpha=0$).
@@ 6.1 The “System 2” Bottleneck
By forcing the token generation to conform to a latent semantic plan, we effectively create a bottleneck of rationality akin to the Global Workspace Theory (baars1988cognitive; bengio2019consciousness; goyal2021coordination). The Idea Head must first commit to a set of concepts, and the Token Head is then constrained to operate within that commitment.
6.2 Efficiency vs. Chain-of-Thought
Recent advancements in reasoning have relied heavily on Chain-of-Thought (CoT) prompting (wei2022chain), which induces planning by generating explicit intermediate tokens. While effective, CoT significantly increases inference latency and cost. Our Idea-Gated architecture offers a parameter-efficient alternative. By encoding the “plan” into a parallel latent vector ($z_{idea}$) rather than a serial token sequence, we achieve semantic control with zero additional decoding steps.
6.3 Limitations
Our experiments highlighted a “Stability-Plasticity Trade-off.” The strong gating mechanism ($\alpha=0.5$) aggressively prunes the search space, which successfully stops drift but increases the risk of repetition loops during greedy decoding. We found that a standard Repetition Penalty ($\rho=1.2$) is necessary to maintain fluency.